How much money would you have if you had 13 quarters, 13 nickels, 13 dimes, and 13 pennies?
Answer by checkley79(3341)

13 quarters 13*25=3.25
13 nickels 13*5=0.65
13 dimes 13*.10=1.00
13 pennies 13*.01=0.13
3.25+0.65+1.00+0.13=$5.03 ans.